OpenStudy (anonymous):

Help me please ! write y= -2/3x + 7 in standar form using integers

OpenStudy (amistre64):

solve it for the form: Ax + By = C

OpenStudy (amistre64):

or maybe some texts have it as: Ax + By + C = 0

OpenStudy (anonymous):

I don't understand, and I need explain to my kid step by step...

OpenStudy (amistre64):

you essentially have: \[y=-\frac abx+c\] sort it out with some algebra if need be into the format:\[ax+by=c\]

OpenStudy (amistre64):

that C is not spose to represent the same number, its just a general constant in either case

OpenStudy (anonymous):

what about the minus?

OpenStudy (anonymous):

sorry I a mom, long time ago out the school :(

OpenStudy (texaschic101):

Standard form is Ax + By = C your equation : y = - 2/3x + 7 you have to rearrange your equation to match the standard form y = - 2/3x + 7 (add 2/3x to both sides) 2/3x + y = -2/3x + 2/3x + 7 (combine like terms) 2/3x + y =7 A = 2/3 B = 1 C = 7 Any questions ?

OpenStudy (anonymous):

thanks thanks a lot, this was complicated for me. I understand better thank you very much. & does the answer have a minus?

OpenStudy (texaschic101):

no.....the answer does not have a minus......if you have anymore questions, please feel free to ask :)

OpenStudy (anonymous):

i do and my answer give 2x + 3y = 21 maybe i have something wrong :(

OpenStudy (texaschic101):

oh...ok....2/3x + y = 7......I forgot, x cannot be a fraction. So, to get rid of the fraction, you have to multiply everything by the denominator of the fraction, which is 3. 2/3x + y = 7 (3)2/3x + (3)y = (3)7 The 3's cancel out leaving you with 2x and 3 *y = 3y and 3*7 = 21 put it all together and you have : 2x + 3y = 21 A = 2 B = 3 C = 21 Sorry about that.....Do you understand ?